# 0pt[Uncaptioned image] Dual loops     0pt[Uncaptioned image] Polycube     0pt[Uncaptioned image] Polycube segmentation

### v3

@@ 4.1 Adding loops
Representation.
Technically speaking, it is possible to represent any possible loop $\ell$ on $\mathcal{M}$ as a function $f\colon S^{1}\rightarrow\mathcal{M}$. However, the continuous search space makes it computationally difficult and expensive to find suitable loops. Instead, we would like to use a discretization of $\mathcal{M}$ to restrict the set of possible loops. A natural choice would be to restrict loops to the vertices and edges of $\mathcal{M}$. However, using this approach, two different loops cannot use the same edge of $\mathcal{M}$. If we would allow two loops to share the same edge, then this could result in ambiguous overlaps between loops and degenerate loop regions. As a result, we may be unable to find any valid loop due to this restriction, making it overly restrictive.
@@
To make it easy to work with this representation of loops, we construct the edge graph $G=(V,E)$ of $\mathcal{M}$. In the edge graph $G$, every node represents an edge of $\mathcal{M}$, and there is an edge between two nodes if the corresponding edges share a triangle in $\mathcal{M}$ (see Figure 4). We also assign a position to the nodes in $G$: if $e=(u,v)$ is an edge of $\mathcal{M}$ (and hence a node in $G$), then we define $p(e)=(p(u)+p(v))/2$. Now, every loop $\ell$ using the representation chosen above simply corresponds to a path in $G$.
Alignment.
A good alignment of the final polycube segmentation can be achieved by creating a polycube map with small distortion. If we consider a loop $\ell$ with axis label $X$, then this corresponds to an $X$-loop on a polycube $Q$. Recall that all points on an $X$-loop have the same $x$-coordinate. Equivalently, all points on an $X$-loop lie in a single plane perpendicular to the $X$-axis. Thus, to achieve good alignment, the loop $\ell$ should have similar properties.
Now consider a point $p$ on loop $\ell$. Let $d(p)$ be the tangent vector at $p$ in the direction of traversal of $\ell$, and let $n(p)$ be the normal vector at $p$ with respect to $\mathcal{M}$. If $\ell$ were an $X$-loop on a polycube, then both $d(p)$ and $n(p)$ would be perpendicular to the $X$-axis. Equivalently, the cross product $d(p)\times n(p)$ would align with the $X$-axis. Therefore, for a good loop $\ell$ with axis label $X$, we would like the angle between $d(p)\times n(p)$ and the $X$-axis to be small for all points $p$ on $\ell$.
@@
We can now compute a loop $\ell$ by first choosing a starting node $e$ in $G$, and then computing the shortest non-empty path from $e$ to itself in $G$ using the weight function $w_{X}$. To enforce the loop to traverse fully around $\mathcal{M}$, we need to make sure that misaligned edges are sufficiently penalized, for otherwise the loop would simply return directly. This can be achieved by setting the strictness factor sufficiently high.
Validity.
A loop $\ell$ computed in $G$ may not be valid, that is, adding $\ell$ to $\mathcal{L}$ may not result in a valid polycube loop structure. The validity of a loop depends on its topological structure. The topological structure of a loop $\ell$ is defined by the sequence of loop regions of $\mathcal{L}$ visited by $\ell$. In [SSV25] it is shown how to efficiently determine if the topological structure of a loop is valid, and how to enumerate all valid (topological structures of) loops that can be added to $\mathcal{L}$.
We now compute a valid loop $\ell$ in $G$ as follows. First, we use the results of [SSV25] to choose a topological structure that corresponds to a valid loop. This topological structure then determines the sequence of loop regions in $\mathcal{L}$ that must be visited by $\ell$. We observe that each loop region can be visited at most once in this sequence. Thus, we can enforce this topological structure in $G$ as follows: for every two consecutive loop regions $R_{i}$ and $R_{i+1}$ in the sequence, we remove all (directed) edges from $G$ that go from $R_{i}$ to another loop region that is not $R_{i+1}$. Now we can use this restricted version of $G$ to compute the most aligned loop $\ell$ that has the chosen topological structure.
To find the best loop possible to add to $\mathcal{L}$, we should try all possible topological structures that correspond to a valid loop. However, as already shown in [SSV25], there may be an exponential number of valid topological structures. We therefore restrict our search to a limited set of topological structures. For each loop region $R$ in $\mathcal{L}$, we consider only the topological structure that visits the fewest loop regions. This approach limits the number of possible valid loops (of some specific axis label) to the number of loop regions in $\mathcal{L}$.
Figure 5: Polycube corners should align with vertices having similar face normals. Patch boundaries should follow mesh edges where adjacent face normals resemble those along the polycube edge.
@@ 4.2 Primalization
Corner placement.
Before we can place corners in loop regions, we first ensure that every loop region contains at least one vertex. To that end, we refine the mesh $\mathcal{M}$ into a mesh $\mathcal{M}^{\prime}$ by adding the loops in $\mathcal{L}$ to $\mathcal{M}$ (properly separated) and by further subdividing loop regions without internal vertices. To find the best vertex $v\in\mathcal{M}^{\prime}$ on which to place a corner, we use two main criteria: (1) The local neighborhood of $v$ in $\mathcal{M}^{\prime}$ should be as similar as possible to the local neighborhood of the corresponding corner in the polycube $Q$ (see Figure 5); (2) Corners that belong to the same zone should be placed on vertices with similar coordinates. Specifically, if two corners belong to the same $X$-zone in the corresponding polycube $Q$, then they should be placed on vertices in $\mathcal{M}^{\prime}$ with a similar $x$-coordinate (see Figure 6).
@@
Figure 6: Polycube corners in the same zone share a coordinate.
@@
Paths.
The next step is to compute non-intersecting paths connecting the corners of adjacent loop regions. We compute these paths incrementally (one after the other), which is both efficient and results in high-quality paths. To give these paths enough freedom to route between corners, we allow them to route on both the edge graph of $\mathcal{M}^{\prime}$ and on the vertices and edges of $\mathcal{M}^{\prime}$ itself. Specifically, we construct a graph $G^{+}$ from $\mathcal{M}^{\prime}$, where every node of $G^{+}$ corresponds to either a vertex or edge of $\mathcal{M}^{\prime}$, and where two nodes of $G^{+}$ are connected by an edge if one of the following holds: (1) the two corresponding vertices are connected by an edge in $\mathcal{M}^{\prime}$, (2) the two corresponding edges share a triangle in $\mathcal{M}^{\prime}$, or (3) the corresponding vertex/edge pair shares a triangle in $\mathcal{M}^{\prime}$.
@@
$a_{l}(e)=\begin{cases}1&\text{if }n_{l}\cdot n(t_{l})\leq 1/\sqrt{3}\\ \text{angle}(n_{l},n(t_{l}))^{2}&\text{otherwise}\end{cases}$
Note that the alignment penalty is $1$ (no penalty) if the angle between $n_{l}$ and $n(t_{l})$ is smaller than $\arccos{1/\sqrt{3}}\approx 0.955$ radians. This is the largest angle between a normal vector and its most similar principal direction. The penalty $a_{r}(e)$ is defined analogously. We now adapt the weight of $e$ by multiplying its current weight (Euclidean distance) by $(a_{l}(e)+a_{r}(e))/2$.
We can now compute a path $P$ between two adjacent corners by finding the shortest path in $G^{+}$ with the given weights. Afterwards, we add the edges of $P$ to the underlying mesh $\mathcal{M}^{\prime}$, and update $\mathcal{M}^{\prime}$ and $G^{+}$ accordingly. We repeat this process until all paths have been computed, after which we obtain a valid polycube segmentation. Note that it is straightforward to assign labels (±$\pm X/\pm Y/\pm Z$) to the triangles/patches of $\mathcal{M}^{\prime}$ given the placement of the corners, the paths, and the loop structure.
@@ 4.3 Loop structure optimization
Initialization.
For models of genus zero, the initial polycube loop structure should correspond to a single cube as polycube. This means that the loop structure has exactly one $X$-loop, one $Y$-loop, and one $Z$-loop. We can use the approach of Section 4.1 to add these $3$ loops in arbitrary order. Note that the topological structures of these initial loops are fixed, although we can still choose in which loop regions we start the search.
For models of higher genus, initialization is significantly more challenging. There are multiple possible corresponding polycubes of minimum complexity, and it is also important to choose the right orientation of the holes in the model. We therefore initialize these loop structures by choosing the topological structures of the initial loops manually. We leave the automated initialization of loop structures for higher-genus models to future work.
Optimization function.
@@
As already discussed at the start of Section 4, we consider two optimization criteria: alignment and complexity. The complexity $f_{c}(\mathcal{L})$ of a loop structure $\mathcal{L}$ is simply measured by the number of loops. To compute the alignment, we first need to primalize the loop structure $\mathcal{L}$. Then, for each triangle $t$ in the resulting mesh $\mathcal{M}^{\prime}$, we consider its normal $n(t)$ and the vector $l(t)$ that corresponds to the label given to $t$ by the polycube segmentation. Let $f_{a}(t)$ be the alignment between $n(t)$ and $l(t)$ computed using the dot product:
$f_{a}(t)=n(t)\cdot l(t)$
@@
Observe that the dot product measures the cosine of the angle between $n(t)$ and $l(t)$. As such, $a(t)$ is a value between $-1$ and $1$, where $1$ indicates perfect alignment.
@@
To compute the total alignment of a polycube segmentation $\mathcal{S}(\mathcal{M}^{\prime})$, we simply use an area-weighted sum:
$f_{a}(\mathcal{S}(\mathcal{M}^{\prime}))=\frac{1}{\text{area}(\mathcal{M}^{\prime})}\sum_{t\in\mathcal{M}^{\prime}}f_{a}(t)*\text{area}(t)$
Observe that the alignment score is always between $-1$ and $1$.
@@
Since we have two optimization criteria, we essentially obtain a multi-objective optimization problem. To solve this problem, we use linear scalarization to turn this problem into a single-objective optimization problem. Specifically, we define the quality of a solution as
$f(\mathcal{L},\mathcal{S}(\mathcal{M}^{\prime}))=f_{a}(\mathcal{S}(\mathcal{M}^{\prime}))-\beta\,f_{c}(\mathcal{L}),$
@@
where $\beta$ is a parameter that controls the trade-off between complexity and alignment. This optimization function can be interpreted as follows. For every additional loop in the loop structure, the total alignment score should increase by at least $\beta$. The parameter $\beta$ can then also be used to control the complexity of the resulting polycube segmentation.
Evolutionary algorithm.
To optimize the polycube loop structure, we use an evolutionary algorithm with only mutation and selection (no crossover). Specifically, our algorithm maintains a population of $N$ loop structures, which are all initialized as described above. In each generation we generate $N^{\prime}$ offspring as follows. First, we pick a loop structure $\mathcal{L}$ from our population uniformly at random. Then we randomly apply one of two possible mutations, each with a $50\%$ chance. In the first mutation we randomly choose a loop $\ell$ to be removed from $\mathcal{L}$, that is, $\mathcal{L}^{\prime}=\mathcal{L}\setminus\{\ell\}$. Here we restrict ourselves to loops $\ell$ such that $\mathcal{L}^{\prime}$ is a valid polycube loop structure (see [SSV25] for how to compute this efficiently). For the second mutation, we choose values $3$ $n_{x}$, $n_{y}$, and $n_{z}$, each uniformly at random in the range $[0,2]$. We then add $n_{x}$ $X$-loops, $n_{y}$ $Y$-loops, and $n_{z}$ $Z$-loops to $\mathcal{L}$ as described in Section 4.1, each starting from a randomly selected position/loop region. Note that all offspring loop structures are valid polycube loop structures by construction.
Given the parent solutions and the offspring solutions, we choose a new generation by selecting the $\lceil N/2\rceil$ best offspring solutions and the $\lfloor N/2\rfloor$ best parent solutions. We continue this process until the best solution in a generation has not improved over the last $10$ generations, at which point we return this best solution, consisting of a loop structure and a corresponding polycube segmentation, as the output of our algorithm.

### v4

@@ 4.1 Adding loops
Representation.  Technically speaking, it is possible to represent any possible loop $\ell$ on $\mathcal{M}$ as a function $f\colon S^{1}\rightarrow\mathcal{M}$. However, the continuous search space makes it computationally difficult and expensive to find suitable loops. Instead, we would like to use a discretization of $\mathcal{M}$ to restrict the set of possible loops. A natural choice would be to restrict loops to the vertices and edges of $\mathcal{M}$. However, using this approach, two different loops cannot use the same edge of $\mathcal{M}$. If we would allow two loops to share the same edge, then this could result in ambiguous overlaps between loops and degenerate loop regions. As a result, we may be unable to find any valid loop due to this restriction, making it overly restrictive.
@@
To make it easy to work with this representation of loops, we construct the edge graph $G$ of $\mathcal{M}$. In the edge graph $G$, every node represents an edge of $\mathcal{M}$, and there is a directed edge (both directions) between two nodes if the corresponding edges share a triangle in $\mathcal{M}$ (see Figure 5). We also assign a position to the nodes in $G$: if $e=(u,v)$ is an edge of $\mathcal{M}$ (and hence a node in $G$), then we define $p(e)=(p(u)+p(v))/2$. Now, every loop $\ell$ using the representation chosen above simply corresponds to a path in $G$.
Alignment.  A good alignment of the final polycube segmentation can be achieved by creating a polycube map with small distortion. If we consider a loop $\ell$ with axis label $X$, then this corresponds to an $X$-loop on a polycube $Q$. Recall that all points on an $X$-loop have the same $x$-coordinate. Equivalently, all points on an $X$-loop lie in a single plane perpendicular to the $X$-axis. Thus, to achieve good alignment, the loop $\ell$ should have similar properties.
Now consider a point $p$ on loop $\ell$. Let $d(p)$ be the tangent vector at $p$ in the direction of traversal of $\ell$, and let $n(p)$ be the normal vector at $p$ with respect to $\mathcal{M}$. If $\ell$ were an $X$-loop on a polycube, then both $d(p)$ and $n(p)$ would be perpendicular to the $X$-axis. The cross product $d(p)\times n(p)$ would align with the $X$-axis. Therefore, for a good loop $\ell$ with axis label $X$, we would like the angle between $d(p)\times n(p)$ and the $X$-axis to be small for all points $p$ on $\ell$.
@@
We can now compute a loop $\ell$ by choosing a starting node $e$ in $G$, and then computing the shortest non-empty path from $e$ to itself in $G$ using the weight function $w_{X}$. To enforce the loop to traverse fully around $\mathcal{M}$, we need to make sure that misaligned edges are sufficiently penalized, otherwise the loop would return instantly. This can be achieved by setting $\alpha$ sufficiently high.
Validity.  Not all loops are valid, that is, adding an arbitrary loop $\ell$ to $\mathcal{L}$ may not result in a valid polycube loop structure. The validity of a loop depends on its topological structure. The topological structure of a loop $\ell$ is defined by the sequence of loop regions of $\mathcal{L}$ visited by $\ell$. In Section 3 we discussed our approach from [SSV25] to enumerate all valid loops.
We compute a valid loop $\ell$ in $G$ as follows. First, we use the graphs $G_{V}^{X}$, $G_{V}^{Y}$, and $G_{V}^{Z}$ from [SSV25] to find a topological structure that corresponds to a valid loop. This topological structure then determines the sequence of loop regions in $\mathcal{L}$ that must be visited by $\ell$. However, there may be an exponential number of valid topological structures. We therefore restrict our search to the following set of topological structures. Specifically, given a starting region $R_{e}$ (which contains the starting edge $e$), we consider the following simple cycles. For all other regions $R^{\prime}\not=R_{e}$, find the simple cycles that travel from $R_{e}$ to $R^{\prime}$ and then return to $R_{e}$, while visiting as few loop regions as possible. This strategy results in at most as many valid topological structures as there are loop regions in $\mathcal{L}$.
We can then compute an actual embedding of the loop for a given topological structure as follows. For each pair of consecutive regions $R_{i}$ and $R_{i+1}$ in the sequence, we remove all directed edges in $G$ that connect $R_{i}$ to any region other than $R_{i+1}$. This restricted version of $G$ ensures that any computed path conforms to the desired topological structure. Finally, within this constrained graph, we compute the most aligned loop $\ell$ that matches the chosen structure. By construction, this loop is guaranteed to be valid.
@@ 4.2 Primalization
Figure 6: Polycube corners should align with vertices having similar face normals.
Figure 7: Polycube corners in the same zone share a coordinate.
Corner placement.  Before placing corners in loop regions, we first ensure that each loop region contains at least one vertex. To this end, we refine the mesh $\mathcal{M}$ into $\mathcal{M}^{\prime}$ using constrained Delaunay triangulation, inserting a vertex into every loop region that does not already contain one. To determine the best vertex $v\in\mathcal{M}^{\prime}$ for placing a corner, we apply two criteria: (1) The local neighborhood of $v$ in $\mathcal{M}^{\prime}$ should be as similar as possible to the local neighborhood of the corresponding corner in the polycube $Q$ (see Figure 6); (2) Corners that belong to the same zone should be placed on vertices with similar coordinates. Specifically, if two corners belong to the same $X$-zone, then they should be placed on vertices in $\mathcal{M}^{\prime}$ with a similar $x$-coordinate (see Figure 7).
Figure 9: Primalization of the polycube loop structure to a polycube segmentation. Notice the placement of the corners and paths.
@@
Paths.  The next step is to compute non-intersecting paths connecting the corners of adjacent loop regions. We compute these paths incrementally (one after the other), which is both efficient and results in high-quality paths. To give these paths enough freedom to route between corners, we allow them to route on both the edge graph of $\mathcal{M}^{\prime}$ and on the vertices and edges of $\mathcal{M}^{\prime}$ itself. Specifically, we construct a graph $G^{+}$ from $\mathcal{M}^{\prime}$, where every node of $G^{+}$ corresponds to either a vertex or edge of $\mathcal{M}^{\prime}$, and where two nodes of $G^{+}$ are connected by an edge if one of the following holds: (1) the two corresponding vertices are connected by an edge in $\mathcal{M}^{\prime}$, (2) the two corresponding edges share a triangle in $\mathcal{M}^{\prime}$, or (3) the corresponding vertex/edge pair shares a triangle in $\mathcal{M}^{\prime}$.
@@
Figure 8: Patch boundaries should follow mesh edges where adjacent face normals resemble those along the polycube edge.
Figure 10: Minimal-complexity polycube configurations for higher genus shapes.
@@
$a_{l}(e)=\begin{cases}1&\text{if }\theta\leq 0.955\\ (1+\theta-0.955)^{2}&\text{otherwise}\\ \end{cases}$
with $\text{with }\theta=\text{angle}(n_{l},n(t_{l}))$
Note that the alignment penalty is $1$ (no penalty) if the angle between $n_{l}$ and $n(t_{l})$ is smaller than $\cos^{-1}(1/\sqrt{3})\approx 0.955$ radians. This is the largest angle between a normal vector and its most similar principal direction. The penalty $a_{r}(e)$ is defined analogously. We now adapt the weight of $e$ by multiplying its current weight (Euclidean distance) by $(a_{l}(e)+a_{r}(e))/2$.
We can now compute a path $P$ between two adjacent corners by finding the shortest path in $G^{+}$ with the given weights. Afterwards, we add the edges of $P$ to the underlying mesh $\mathcal{M}^{\prime}$, and update $\mathcal{M}^{\prime}$ and $G^{+}$ accordingly. We repeat this process until all paths have been computed, after which we obtain a valid polycube segmentation. Note that it is straightforward to assign labels (±$\pm X/\pm Y/\pm Z$) to the triangles/patches of $\mathcal{M}^{\prime}$ given the placement of the corners, the paths, and the loop structure, see Figure 9.
@@ 4.3 Loop structure optimization
Initialization.  For genus-zero models, the initial polycube loop structure should correspond to a single cube. This requires exactly one $X$-loop, one $Y$-loop, and one $Z$-loop. Using the approach from Section 4.1, we can add these loops in any order. Although their topological structures are fixed, we can still vary the starting loop region for each. To obtain a good starting configuration, we generate 100 candidate solutions as follows: randomly choose an order for the $X$, $Y$, and $Z$ loops. For each loop, sample a random edge as the starting point. The first loop can be any valid loop. The second must intersect the first loop twice. The third must intersect both the first and second loops twice, in alternating fashion. This guarantees a valid initial loop structure. Among the 100 samples, we initialize with the one that yields the best polycube segmentation according to the criteria in Section 4.
For models with higher genus, initialization becomes significantly more challenging. Unlike the genus-zero case, which always starts with a single cube, higher-genus shapes admit many distinct minimal-complexity polycube configurations. This is because each handle introduces degrees of freedom in its orientation. For example, a polycube torus can be oriented in three different ways, see Figure 10(a,b,c). A genus-2 surface has two independent holes, each of which can be oriented separately, further increasing the number of valid configurations, see Figure 10(d,e). There is currently no robust method to automatically determine an optimal initial polycube or to embed the corresponding loops on the surface. To address this, we initialize the loop structures manually. The user selects starting edges on the mesh, and for each selected edge, the method from Section 4.1 computes a valid loop. By iteratively adding such loops, the user builds a minimal-complexity polycube loop structure. This process is relatively straightforward for users familiar with the appearance of higher-genus polycube structures on the surface. For genus up to 5, this manual initialization typically takes less than 10 minutes. For significantly higher genus, however, manual initialization becomes more difficult. Automating this process remains an open problem for future work.
(a) $\beta=0$
(b) $\beta=0.001$
(c) $\beta=0.01$
(d) $\beta=0.1$
Figure 11: Influence of $\beta$ on compactness of resulting polycube.
@@
Optimization function.  As already discussed at the start of Section 4, we consider two optimization criteria: fidelity and compactness. The compactness $C(\mathcal{L})$ of a loop structure $\mathcal{L}$ is simply measured by the number of loops. To compute the fidelity, we first need to primalize the loop structure $\mathcal{L}$. Then, for each triangle $t$ in the resulting mesh $\mathcal{M}^{\prime}$, we consider its normal $n(t)$ and the vector $l(t)$ that corresponds to the label given to $t$ by the polycube segmentation. Let $f(t)$ be the alignment between $n(t)$ and $l(t)$ computed using the dot product:
$f(t)=n(t)\cdot l(t)$
@@
Observe that the dot product measures the cosine of the angle between $n(t)$ and $l(t)$. As such, $f(t)$ is a value between $-1$ and $1$, where $1$ indicates perfect alignment.
@@
To compute the total fidelity $F(\mathcal{S}(\mathcal{M}^{\prime}))$ of a polycube segmentation $\mathcal{S}(\mathcal{M}^{\prime})$, we simply use an area-weighted sum:
$F(\mathcal{S}(\mathcal{M}^{\prime}))=\frac{1}{\text{area}(\mathcal{M}^{\prime})}\sum_{t\in\mathcal{M}^{\prime}}f(t)*\text{area}(t)$
Observe that the value for fidelity is always between $-1$ and $1$.
@@
Since we have two optimization criteria, we essentially obtain a multi-objective optimization problem. To solve this problem, we use linear scalarization to turn this problem into a single-objective optimization (maximization) problem. Specifically, we define the quality of a solution as
$\text{quality}(\mathcal{L},\mathcal{S}(\mathcal{M}^{\prime}))=F(\mathcal{S}(\mathcal{M}^{\prime}))-\beta*C(\mathcal{L}),$
@@
where $\beta$ is a parameter that controls the trade-off between compactness and fidelity. This optimization function can be interpreted as follows. For every additional loop in the loop structure, the total fidelity score should increase by at least $\beta$. The parameter $\beta$ can then also be used to control the compactness of the resulting polycube, see Figure 11.
Evolutionary algorithm.  We optimize the polycube loop structure using an evolutionary algorithm with mutation and selection only (no crossover). The algorithm maintains a population of $\mu$ loop structures, initialized as described before. In each generation, $\lambda$ offspring are produced by mutating randomly selected parents.
For each offspring, we draw a random value in $[0,1]$; if it exceeds $\tau$, we apply an addition mutation, otherwise a removal mutation. In the addition mutation, we sample $n_{x},n_{y},n_{z}\in[\delta_{\min},\delta_{\max}]$ and insert the corresponding number of $X$-, $Y$-, and $Z$-loops at random positions (Section 4.1). Note that for this we must compute the $G_{V}^{X}$, $G_{V}^{Y}$, $G_{V}^{Z}$ graphs first. Then for each loop to be added, we first compute a set of valid topological structures, and grab a random valid topological structure to be embedded on $\mathcal{M}$. In the removal mutation, we sample $n_{r}\in[\delta^{\prime}_{\min},\delta^{\prime}_{\max}]$ and delete that many valid loops (Section 3). All mutations preserve validity.
The next generation consists of the top $\lceil\mu/2\rceil$ offspring and $\lfloor\mu/2\rfloor$ parents, selected by the optimization criteria explained above. The process terminates after $\chi$ generations without improvement, returning the best solution found.
We use $\mu=10$, $\lambda=30$, $\tau=0.5$, $\delta_{\min}=0$, $\delta_{\max}=2$, $\delta^{\prime}_{\min}=1$, $\delta^{\prime}_{\max}=1$, and $\chi=10$, which we found to yield satisfactory results across a range of input meshes.
Input: Input triangle mesh $\mathcal{M}=(V,T)$
Output: Polycube segmentation $S(\mathcal{M})$
1
[…]
3
[…]
6
[…]
8       $\mathcal{O}\leftarrow\emptyset$
9
10      for $i\leftarrow 1$ to $\lambda$ do
11             $L\leftarrow$ random element from $\mathcal{P}$
12
13            if random value $\in[0,1]$ $>\tau$ then
14                   $n_{X},n_{Y},n_{Z}\leftarrow$ random integers in $[\delta_{\min},\delta_{\max}]$
15                   Add $n_{X}$ $X$-loops, $n_{Y}$ $Y$-loops, and $n_{Z}$ $Z$-loops to $L$
16
17            else
18                   $n_{r}\leftarrow$ random integer in $[\delta^{\prime}_{\text{min}},\delta^{\prime}_{\text{max}}]$
19                   Remove $n_{r}$ loops from $L$
20
21
22            Add $L$ to $\mathcal{O}$
23
24
[…]
26
[…]
32
33
Algorithm 1 DualCube
4.4 Analysis
We analyze the computational complexity of our method using $n^{\prime}$ to denote the number of vertices in the refined mesh $\mathcal{M}^{\prime}$. Since $\mathcal{M}^{\prime}$ is a refinement of a triangular mesh, the number of faces and edges is $O(n^{\prime})$. The edge graph $G$ also has $O(n^{\prime})$ elements, as it contains a constant number of elements per element in $\mathcal{M}^{\prime}$. Let $l$, $s$, $r$, and $z$ denote the number of loops, loop segments, loop regions, and zones, respectively. Note that, if we have $r$ loop regions, then we can have at most $6r$ loop segments. Furthermore, if we have $6r$ loop segments, then we can have at most $6r$ loops. Lastly, if we have $r$ loop regions, then we have at most $3r$ zones.
Loop addition.  To add a loop, we construct graph $G_{L}$ with $s$ vertices and at most $6s$ edges. We filter $G_{L}$ to obtain the valid graphs $G_{V}^{X}$, $G_{V}^{Y}$, and $G_{V}^{Z}$ in $O(s)$ time. We run $r$ modified Dijkstra runs on $G_{V}^{X}$, $G_{V}^{Y}$, or $G_{V}^{Z}$ to find at most $r$ simple cycles in $O(rs\log s)$. We embed one of the loops on $\mathcal{M}^{\prime}$ with one Dijkstra run on edge graph $G$, which takes $O(n^{\prime}\log n^{\prime})$ time. Thus, each addition costs $O(rs\log s+n^{\prime}\log n^{\prime})$ time, or $O(r^{2}\log r+n^{\prime}\log n^{\prime})$ time.
Loop removal.  Loop removal checks Conditions 1–4 in $O(s)$ time. Condition 5 is checked via its level graph with $z$ nodes and $l$ edges. We detect cycles in these graphs using DFS in $O(z+l)$ time. Thus, each removal costs $O(s+z+l)$ time, or $O(r)$ time.
Primalization.  Constrained Delaunay refinement on $\mathcal{M}^{\prime}$ takes $O(n^{\prime}\log n^{\prime})$ time. Assigning naive labels to faces and computing local label sets for vertices takes $O(n^{\prime})$. We compute label similarity in $O(n^{\prime})$ time and sort the vertices by similarity in $O(n^{\prime}\log n^{\prime})$ time. For each zone, we sort candidates three times by $x$, $y$, and $z$ coordinates in $O(n^{\prime}\log n^{\prime})$ time. For each loop region, we choose the best candidate in $O(n^{\prime})$ time. For each loop segment, we compute a path via Dijkstra on $\mathcal{M}^{\prime}$ in $O(sn^{\prime}\log n^{\prime})$ time. Thus, primalization costs $O(sn^{\prime}\log n^{\prime})$ time, or $O(rn^{\prime}\log n^{\prime})$ time. Fidelity and compactness can be calculated in $O(n^{\prime}+l)$ time, or $O(n^{\prime}+r)$ time.
Automatic initialization involves constructing $\mu$ initial loop structures using 3 loop additions in $O(r^{2}\log r+n^{\prime}\log n^{\prime})$ time. In each iteration, each offspring requires at most $3\delta_{\text{max}}$ loop additions, $\delta^{\prime}_{\text{max}}$ loop removals, and a primalization step. Thus, the cost per individual per iteration is $O(\delta_{\text{max}}(r^{2}\log r+n^{\prime}\log n^{\prime})+\delta^{\prime}_{\text{max}}r)$.
Expressing this in terms of the input mesh vertex count $n$, we observe that $n^{\prime}\leq n+r$ (each loop region introduces at most one new vertex). Furthermore, the number of loop regions $r$ corresponds to the number of polycube corners in the final solution. It is fair to assume that the resulting polycube contains significantly fewer corners than the number of vertices in the input mesh $\mathcal{M}$, i.e., $r\ll n$. Under these assumptions, we get that the runtimes simplify to:
initialization per individual $O(n^{2}\log n)$
per iteration per individual $O(\delta_{\text{max}}(n^{2}\log n)+\delta^{\prime}_{\text{max}}n)$